OpenStudy (anonymous):

Find the derivative for y=e^-2xcos(6x)

OpenStudy (campbell_st):

you need to product rule. where \[h(x) = e^{-2x}\] and \[g(x) = \cos(6x)\] so \[\frac{dy}{dx} = g(x)\times h'(x) + h(x) \times g'(x)\]

OpenStudy (anonymous):

Ok, so let me know if I'm on the right track. dy/dx= cos(6x) * -2xe + e^-2x * 6xcos

OpenStudy (campbell_st):

that looks good except for the derivative of cos(6x) dy/dx = -6sin(6x)